a cone with a height of 10 in and a baseof radius 3 in
Answer by checkley77(12844)

V=BH/3
V=PIR^2*10/3
V=3.14*3^2*10/3
V=3.14*9*10/3
V=3.14*90/3
V=3.14*30
V=94.2 IN^3 ANS.